OpenStudy (anonymous):

Find the principal root. Express the result in the form a + bi with a and b rounded to the nearest hundredth. (-1)^(1/4)

OpenStudy (anonymous):

@Nnesha @jdoe0001

OpenStudy (jdoe0001):

I assume you don't have any choices?

OpenStudy (anonymous):

sorry but no.

OpenStudy (jdoe0001):

hmm all what I can see is \(\large \bf (-1)^{\frac{1}{4}}\implies (-1^{\frac{1}{2}})^{\frac{1}{2}}\implies \sqrt{\quad \sqrt{-1}\quad }\implies \sqrt{i} \)

OpenStudy (jdoe0001):

from which I do get a value from my calculator, though dunno where it gets it from =)

OpenStudy (anonymous):

ok, thnx

OpenStudy (jdoe0001):

yw